# The Born rule as a natural transformation of functors

### v1

@@ 1 Introduction
Category theory makes mathematically precise the concepts of being and becoming, which perhaps are the most fundamental concepts in all of physics. In particular, a category consists of a collection of objects—which may be thought of as all possible states of being associated with a particular type of structural entity—and a collection of morphisms—which may be thought of as a class of admissible transformations between the various states of being. While category theory was first introduced in the context of pure mathematics [6, 12], it has since been utilized across a wide number of disciplines, including computer science, linguistics, neuroscience, and philosophy. In the context of quantum physics, category theory is central to the construction of topological quantum field theories [2], homological mirror symmetry in string theory [11], topos-theoretic approaches to quantum foundations [5, 10, 17], a systematic study of Feynman diagrams [3], conformal field theory [18, 16], the mysterious connection between path integrals and multiple zeta-values [13], quantum Bayesian and statistical inference [14, 15, 19, 7], and a diagrammatic formulation of quantum information-theoretic protocols [1, 8].
@@
Despite the amount of interest in category theory in the context of quantum physics, it seems to have been overlooked that the basic constituents of quantum theory itself—namely, the notions of quantum state, measurement, and the Born rule—are all simultaneously encoded by the categorical notion of a natural transformation of functors, as we show in this work. A functor is a mapping between categories that respects all of the categorical structure, i.e., it sends objects to objects and morphisms to morphisms. As such, a functor formalizes the notion of analogy, as it translates statements in one category to analogous statements in another category. A natural transformation then formalizes the notion of a mapping between functors, and is arguably the most fundamental notion of category theory. Here we consider the measurement and probability functors—which capture the way in which POVM elements and probabilities are additive with respect to a coarse-graining of measurements—and prove that a quantum state uniquely determines a natural transformation between the measurement and probability functors via the Born rule. Moreover, we prove that every such natural transformation between the measurement and probability functors is induced by the Born rule associated with a unique quantum state.
@@
In more precise mathematical terms, we prove Theorem 5.3, which states that given a fixed quantum system $A$, there exists a bijection from the set of density operators on the associated Hilbert space $\mathcal{H}_{A}$ to the set of natural transformations between the measurement and probability functors. To prove surjectivity, we make use of the Busch-Gleason Theorem [9, 4], which states that every generalized probability measure on the space of effects (i.e., POVM elements on $A$) is induced by the Born rule associated with a unique density operator $\rho$ on $\mathcal{H}_{A}$. As such, there is a precise sense in which Theorem 5.3 may be viewed as a categorical lift of the Busch-Gleason Theorem.
@@
In what follows, we recall the basic notions needed to define a natural transformation of functors in Section 2, and we give some illustrative examples. In particular, we go over the basic construction of the category $\mathfrak{Meas}$ consisting of measurable functions between measurable spaces, which plays a fundamental role in this work. In Section 3 we introduce the measurement and probability functors—which are functors from the category $\mathfrak{Meas}$ to the category $\mathfrak{Set}$ of functions between sets—and show how such functors capture the fact that quantum effects and their associated probabilities are additive with respect to a coarse-graining of measurements. In Section 4, we prove that a natural transformation between the measurement and probably functors uniquely determines a generalized probability measure on the space of effects of a quantum system $A$, thus establishing a direct connection with the Busch-Gleason Theorem. In Section 5, we show how a density operator uniquely determines a natural transformation from the measurement functor to the probability functor via the Born rule, thus setting the stage for the proof of our main result Theorem 5.3.
@@ 2 Categories, functors, and natural transformations
  • a function $\mathbb{dom}:\text{mor}(\mathscr{C})\to\text{ob}(\mathscr{C})$.
  • a function $\mathbb{cod}:\text{mor}(\mathscr{C})\to\text{ob}(\mathscr{C})$.
  • for every triple of objects $(X,Y,Z)$, a binary operation $\text{Hom}(X,Y)\times\text{Hom}(Y,Z)\to\text{Hom}(X,Z)$, which is referred to as composition of morphisms. Here, $\text{Hom}(X,Y)$ denotes the subclass of morphisms $f:X\to Y$ in $\text{mor}(\mathscr{C})$ such that $\text{dom}(f)=X$ and $\text{cod}(f)=Y$.
@@
A measurable space consists of a pair $(X,\Sigma_{X})$, where $X$ is a set and $\Sigma_{X}$ is a collection of subsets of $X$ which are referred to as events. The space of events $\Sigma_{X}$ is required to satisfy the following properties:
@@
  • $\{E_{n}\}_{n=1}^{\infty}\subset\Sigma_{X}\implies\bigcup_{n=1}^{\infty}E_{n}\in\Sigma_{X}$.
@@
A natural transformation from a functor $\mathbb{F}:\mathscr{C}\to\mathscr{D}$ to a functor $\mathbb{G}:\mathscr{C}\to\mathscr{D}$ is a mapping $\mathcal{N}:\mathbb{F}\to\mathbb{G}$ that associates every $X\in\text{ob}(\mathscr{C})$ with a morphism $\mathcal{N}_{X}:\mathbb{F}(X)\to\mathbb{G}(X)$ in $\mathscr{D}$, such that for every morphism $f:X\to Y$ in $\mathscr{C}$, the following diagram commutes.
${\mathbb{F}(X))}$${\mathbb{F}(Y)}$${\mathbb{G}(X)}$${\mathbb{G}(Y)}$$\scriptstyle{\mathbb{F}(f)}$$\scriptstyle{\mathcal{N}_{X}}$$\scriptstyle{\mathcal{N}_{Y}}$$\scriptstyle{\mathbb{F}(g)}$
@@ 3 The measurement and probability functors
In this section, we introduce the measurement and probability functors, which are functors from the category $\mathfrak{Meas}$ consisting of measurable functions between measurable spaces, to the category $\mathfrak{Set}$ consisting of functions between sets. To define the measurement and probability functors, we first need to set some notation and terminology. Throughout this work, we let $A$ denote a quantum system with Hilbert space $\mathcal{H}_{A}$, and the vector space of self-adjoint operators on $\mathcal{H}_{A}$ will be denoted by $\mathbb{Obs}(A)$. The identity operator on $\mathcal{H}_{A}$ will be denoted by $\mathds{1}$ and the set of density operators on $\mathcal{H}_{A}$ will be denoted by $\mathfrak{D}(A)$. Given a measurable space $(X,\Sigma_{X})$, a function $\mu:\Sigma_{X}\to\mathbb{Obs}(A)$ is said to be a positive operator-valued measure (POVM) if and only if $\mu$ satisfies the following properties:
@@
  • $\mu\left(\bigsqcup_{n=1}^{\infty}E_{n}\right)=\sum_{n=1}^{\infty}\mu(E_{n})$.
@@
While it is well known to mathematicians that the measurement and probability functors are indeed functors, we nevertheless give a proof of this fact for the sake of being self-contained.
@@
$\Big{[}\mathbb{M}(\mathrm{id}_{X})(\mu)\Big{]}(E)={\mathrm{id}_{X}}_{*}\mu=\mu(\mathrm{id}^{-1}(E))=\mu(E)\,.$
@@ 5 Quantum states as natural transformations
Injectivity. Let $\rho,\sigma\in\mathfrak{D}(A)$ be two distinct density operators, and let $\Delta\in\mathbb{Obs}(A)$ be the self-adjoint operator given by $\Delta=\rho-\sigma\neq 0$. Clearly ${\rm Tr}[\Delta]=9$ as density operators are of unit trace. By the spectral theorem for self-adjoint operators, $\Delta$ may be written in terms of the spectral integral

### v2

@@ 1 Introduction
Category theory makes mathematically precise the concepts of being and becoming, which perhaps are the most fundamental concepts in all of physics. In particular, a category consists of a collection of objects—which may be thought of as all possible states of being associated with a particular type of structural entity—and a collection of morphisms—which may be thought of as a class of admissible transformations between the various states of being. While category theory was first introduced in the context of pure mathematics [1, 2], it has since been utilized across a wide number of disciplines, including computer science, linguistics, neuroscience, and philosophy. In the context of quantum physics, category theory is central to the construction of topological quantum field theories [3], homological mirror symmetry in string theory [4], topos-theoretic approaches to quantum foundations [5, 6, 7], a systematic study of Feynman diagrams [8], conformal field theory [9, 10], higher gauge theory and quantum gravity [11, 12, 13], the mysterious connection between path integrals and multiple zeta-values [14], quantum Bayesian inference and retrodiction [15, 16, 17, 18, 19], and a diagrammatic formulation of quantum information-theoretic protocols [20, 21].
@@
Despite the amount of interest in category theory in the context of quantum physics, it seems to have been overlooked that the basic constituents of quantum theory itself—namely, the notions of quantum state, measurement, and the Born rule—are all simultaneously encoded by the categorical notion of a natural transformation of functors, as we show in this work. A functor is a mapping between categories that respects all of the categorical structure, i.e., it sends objects to objects and morphisms to morphisms. As such, a functor formalizes the notion of analogy, as it translates statements in one category to analogous statements in another category. A natural transformation then formalizes the notion of a mapping between functors, and is arguably the most fundamental notion of category theory.
Here we consider the measurement and probability functors—which capture the way in which POVM elements and probabilities are additive with respect to a coarse-graining of measurements—and prove that a quantum state uniquely determines a natural transformation between the measurement and probability functors via the Born rule. Moreover, we prove that every such natural transformation between the measurement and probability functors is induced by the Born rule associated with a unique quantum state.
@@
In more precise mathematical terms, we prove Theorem 5.3, which states that given a fixed quantum system $A$, there exists a bijection from the set of density operators on the associated Hilbert space $\mathcal{H}_{A}$ to the set of natural transformations between the measurement and probability functors. To prove surjectivity, we make use of the Busch-Gleason Theorem [22, 23], which states that every generalized probability measure on the space of effects (i.e., POVM elements on $A$) is induced by the Born rule associated with a unique density operator $\rho$ on $\mathcal{H}_{A}$. Thus, there is a precise sense in which Theorem 5.3 may be viewed as a categorical lift of the Busch-Gleason Theorem.
As a consequence of Theorem 5.3, it follows that once a Hilbert space $\mathcal{H}_{A}$ is associated with a quantum system $A$, the traditional separate postulates for quantum states, measurement, and the Born rule become redundant. In particular, these concepts are all subsumed within the single notion of a natural transformation between the measurement and probability functors associated with $\mathcal{H}_{A}$. Furthermore, such natural transformations inherently describe the dynamics of quantum effects and their associated probabilities in relation to the coarse-graining of measurements—a central component in the emerging theories of process matrices/tensors [24, 25, 26] and quantum combs [27]. As such, our results provide yet another example of category theory’s power in unifying core physical concepts.
@@
In what follows, we recall the basic notions needed to define a natural transformation of functors in Section 2, and we give some illustrative examples. In particular, we go over the basic construction of the category $\mathfrak{Meas}$ consisting of measurable functions between measurable spaces, which plays a fundamental role in this work. In Section 3 we introduce the measurement and probability functors—which are functors from the category $\mathfrak{Meas}$ to the category $\mathfrak{Set}$ of functions between sets—and show how such functors capture the fact that quantum effects and their associated probabilities are additive with respect to a coarse-graining of measurements. In Section 4, we prove that a natural transformation between the measurement and probability functors uniquely determines a generalized probability measure on the space of effects of a quantum system $A$, thus establishing a direct connection with the Busch-Gleason Theorem. In Section 5, we show how a density operator uniquely determines a natural transformation from the measurement functor to the probability functor via the Born rule, thus setting the stage for the proof of our main result Theorem 5.3.
@@ 2 Categories, functors, and natural transformations
  • a function $\mathbb{dom}:\text{mor}(\mathscr{C})\to\text{ob}(\mathscr{C})$ given by $\mathbb{dom}(X\to Y)=X$ for every morphism $X\to Y$ in $\mathscr{C}$.
  • a function $\mathbb{cod}:\text{mor}(\mathscr{C})\to\text{ob}(\mathscr{C})$ given by $\mathbb{cod}(X\to Y)=Y$ for every morphism $X\to Y$ in $\mathscr{C}$.
  • for every triple of objects $(X,Y,Z)$, a binary operation $\text{Hom}(X,Y)\times\text{Hom}(Y,Z)\to\text{Hom}(X,Z)$, which is referred to as composition of morphisms. Here, $\text{Hom}(X,Y)$ denotes the subclass of morphisms $f:X\to Y$ in $\text{mor}(\mathscr{C})$ such that $\text{dom}(f)=X$ and $\text{cod}(f)=Y$, and similarly for $\text{Hom}(Y,Z)$ and $\text{Hom}(X,Z)$. Given $(f,g)\in\text{Hom}(X,Y)\times\text{Hom}(Y,Z)$, the composition of $f$ and $g$ will be denoted by $g\circ f$.
@@
A measurable space consists of a pair $(X,\Sigma_{X})$, where $X$ is a set and $\Sigma_{X}$ is a $\sigma$-algebra of subsets of $X$, whose elements will be referred to as events. As the space of events $\Sigma_{X}$ is a $\sigma$-algebra, it is required to satisfy the following properties:
@@
  • If $\{E_{n}\}_{n=1}^{\infty}$ is a countable collection of sets in $\Sigma_{X}$, then their union $\bigcup_{n=1}^{\infty}E_{n}\in\Sigma_{X}$.
@@
A natural transformation from a functor $\mathbb{F}:\mathscr{C}\to\mathscr{D}$ to a functor $\mathbb{G}:\mathscr{C}\to\mathscr{D}$ is a mapping $\mathcal{N}:\mathbb{F}\to\mathbb{G}$ that associates every $X\in\text{ob}(\mathscr{C})$ with a morphism $\mathcal{N}_{X}:\mathbb{F}(X)\to\mathbb{G}(X)$ in $\mathscr{D}$, such that for every morphism $f:X\to Y$ in $\mathscr{C}$ we have $\mathbb{G}(f)\circ\mathcal{N}_{X}=\mathcal{N}_{Y}\circ\mathbb{F}(f)$, i.e., the following diagram commutes.
${\mathbb{F}(X)}$${\mathbb{F}(Y)}$${\mathbb{G}(X)}$${\mathbb{G}(Y)}$$\scriptstyle{\mathbb{F}(f)}$$\scriptstyle{\mathcal{N}_{X}}$$\scriptstyle{\mathcal{N}_{Y}}$$\scriptstyle{\mathbb{G}(f)}$
@@ 3 The measurement and probability functors
In this section, we introduce the measurement and probability functors, which are functors from the category $\mathfrak{Meas}$ consisting of measurable functions between measurable spaces, to the category $\mathfrak{Set}$ consisting of functions between sets. To define the measurement and probability functors, we first need to set some notation and terminology. Throughout this work, we let $A$ denote a quantum system with a (separable) Hilbert space $\mathcal{H}_{A}$, which may be infinite-dimensional. The vector space of bounded self-adjoint operators on $\mathcal{H}_{A}$ will be denoted by $\mathbb{Obs}(A)$. The identity operator on $\mathcal{H}_{A}$ will be denoted by $\mathds{1}$, and the set of density operators—i.e., the set of positive, trace-class operators on $\mathcal{H}_{A}$ of unit trace—will be denoted by $\mathfrak{D}(A)$. We note that for finite-dimensional systems, a density operator is simply a Hermitian operator on $\mathcal{H}_{A}$ whose finite set of eigenvalues are all non-negative and sum to 1. Given a measurable space $(X,\Sigma_{X})$, a function $\mu:\Sigma_{X}\to\mathbb{Obs}(A)$ is said to be a positive operator-valued measure (POVM) if and only if $\mu$ satisfies the following properties:
@@
  • $\mu\left(\bigsqcup_{n=1}^{\infty}E_{n}\right)=\sum_{n=1}^{\infty}\mu(E_{n})$ for every sequence of disjoint events $\{E_{n}\}$, where convergence is with respect to the ultra-weak operator topology on $\mathbb{Obs}(A)$.
@@
While it is well known to mathematicians that the measurement and probability functors are indeed functors, we nevertheless give a proof of this fact for the more general reader.
@@
$\Big[\mathbb{M}(\mathrm{id}_{X})(\mu)\Big](E)=({\mathrm{id}_{X}}_{*}\mu)(E)=\mu(\mathrm{id}^{-1}(E))=\mu(E)\,.$
@@ 5 Quantum states as natural transformations
Injectivity. Let $\rho,\sigma\in\mathfrak{D}(A)$ be two distinct density operators, and let $\Delta\in\mathbb{Obs}(A)$ be the self-adjoint operator given by $\Delta=\rho-\sigma\neq 0$. Clearly ${\rm Tr}[\Delta]=0$ as density operators are of unit trace. By the spectral theorem for self-adjoint operators, $\Delta$ may be written in terms of the spectral integral